OpenStudy (anonymous):

Solve cos x tan x -sin^2x=0 for all real values of x

OpenStudy (anonymous):

tan=sin/cos cos*tan=sin sinx-sin^2x=0 let u=sinx u-u^2=0 hence u=0,u=1 u=0 => sin(x) =0 hence x=k*pi u=1=> sin(x)=1 hence x=pi/2+2kp for any whole number k(that is k is a member of Z) ex. -3,-2,-1,0,1,2,3

OpenStudy (anonymous):

Solve 2 sin^2x-sinx=0 for principal values x 60 and 120 degrees 0 and 150 degrees 0 and 30 degrees 60 degrees